The ages of three brothers are consecutive even integers. Three times the age of the youngest brother exceeds the oldest brothers age by 48 years. What's the age of the youngest brother?

Answer by Boreal(15235)

ages are x, x+2, x+4
3(x)-48=x+4 is the equation or 3(x)=x+4+48
2x=52
x=26 ANSWER
the other two are 28 and 30
3(26)-48=30.